M ICH I Water for tetrachloroethene and led hloroethane assuming water solubilities of 150 and 8 mg/L, respectively. Also, sk solve for #3
Answer #1

3) a)

Given : 1) Mass of methane ( m) = 24.0 g

2) volume of vessel = 5.00 L

3) temperature of gas = 0 0 C = 273.15 K

We have relation , P V = n R T

where, P is a pressure of a gas, V is a volume of a gas, n is no. of moles of gas, R is a gas constant and T is temperature of gas.

\therefore P = n R T / V

We know that, no. of moles = Mass / molar mass

Molar mass of CH 4 = 12.01 + ( 4 \times 1.0079 ) = 16.04 g/ mol

Hence, no. of moles of methane gas = 24.0 g /16.04 g/mol = 1.496 mol

\therefore Pressure exerted by methane gas = 1.496 mol \times 0.082057 L atm mol -1 K -1\times 273.15 K / 5.00 L

Pressure exerted by methane gas = 6.706 atm

ANSWER : Pressure exerted by methane gas = 6.71 atm

b)

We have relation, P V = n R T

But n = mass ( m ) / Molar mass (M)

\therefore P V = (m / M ) R T

On rearranging above equation , we get

P M = (m / V ) R T

\therefore P M = d R T

\therefore density of gas = P M / R T

Given : Temperature = 20 + 273 .15 = 293.15 K

Pressure of gas = 0.967 bar = 0.9543 atm

Molar mass of methane gas = 16.04 g /mol

\therefore density of methane gas = 0.9543 atm \times 16.04 g /mol / 0.082057 L atm mol -1 K -1\times 293.15 K

= 0.636 g / L

= 0.636 g / 1000 ml

= 0.000636 g / ml

ANSWER : Density of methane gas = 0.000636 g /ml